# Multispectral airborne laser scanning for tree species classification: a benchmark of machine learning and deep learning algorithms

### v1

@@ 2.3 Pre-processing of laser scanning data
The data from each scanner were loaded to TerraScan (Terrasolid Ltd., Finland), and for each point the closest points within 20 cm were sought from the other two channels to form the reflectance triplets that were then stored in the RGB color fields in the LAZ file structure. Thus, points that do not have close-by counterparts on either one or two channels may exhibit only NA (coded as value ’1’) on the respective RGB fields. The reflectance values in the LAZ were stored as decibel values and exported with RIEGL extrabytes (echo deviation, reflectance, and pulse width). Eventually, the full point cloud dataset was split into 200 m square tiles and delivered for the benchmark in 1.2 LAZ format with ETRS-TM35FIN coordinates and ellipsoidal elevation values.
Preprocessing steps for the Optech Titan data included some basic processing steps carried out by TerraTec Oy, intensity range (R) correction according to $\mathrm{R}^{2}$ (Matikainen et al., 2017), cutting of overlap points and removal of an additional cross line (using TerraScan). The point densities measured after preprocessing were 11 points/$\mathrm{m}^{2}$, 13 points/$\mathrm{m}^{2}$ and 11 points/$\mathrm{m}^{2}$ for the Channels 1, 2 and 3, respectively. The preprocessed set of data was provided as ‘original’ data for the benchmarking.
@@
Tree segmentation was carried out based on the Optech Titan data. The resulting segments were also applied for tree species classification with the HeliALS data to eliminate the variability of segment boundaries from the comparison. Given the fact that plenty of deciduous trees are growing in the study area, we used a varying window watershed-based method for the tree detection first depicted in (Kaartinen et al., 2012). First, we created a canopy height model (CHM) from the first returns of the vegetation points of Channels 1 and 2 with 0.5 m pixel size. The CHM was then smoothed by a Gaussian filter with a varying window size, followed by a local maximum filtering to identify the tops of trees. The window size varied according to the CHM height of the target as follows: 3 pixels for 0–7 m height, 5 pixels for 7–20 m height, 7 pixels for 20–30 m height, and 9 pixels for greater than 30 m height. Finally, the marker-controlled watershed algorithm was applied to delineate the crowns of trees using detected tree tops as seeds (Kaartinen et al., 2012). Output of the tree segmentation is a raster image with a unique label for each segment, i.e. the pixels with the same label form one tree. These images together with the original and normalized point clouds were delivered to the partners for species classification.
@@ 2.4 Benchmark dataset
Figure 3: (a) Example multispectral point clouds and photographs of the nine different tree species found in the dataset. The red, green and blue color channels of the point clouds represent the laser pulse return intensities at the wavelengths $\lambda_{1}=\text{532 nm},\lambda_{2}=\text{905 nm}$ and $\lambda_{3}=\text{1550 nm}$, respectively. The height scale varies between the visualized point clouds. Photographs of the representative species instances were taken during the autumn of 2024. (b) Number of segments per species in the training and test sets. (c) Species-wise tree height distributions for the entire dataset.
The contestants of the benchmark competition received a training set containing the segment labels and the ground-truth species classes. The training set consisted of 1065 segments. The participating methods were benchmarked with a test set containing 5261 segments. Figure 3 visualizes the nine tree species included in the dataset, the number of segments per species in the training and test sets, and also the species-wise tree height distributions for the entire dataset.
Furthermore, the entire dataset was used to study the scaling of the classification error as a function of the training set size and the point density. We conducted these experiments using the deep learning and machine learning models with the highest overall accuracy on the HeliALS dataset in the benchmark competition. In the scaling law studies, we estimated the classification errors of the two models using 5-fold cross-validation. The folds were selected based on stratified random sampling to ensure that each fold follows the same species distribution. In the scaling law experiments with a varying number of training segments, we created training sets ranging from 250 segments to 5000 segments (in 250 segment increments) with a similar species distribution using another round of stratified random sub-sampling from the folds previously generated from the full dataset.
@@ 2.5 Reference data collection
A browser-based crowdsourcing tool was developed to help collecting the ground-truth species data in the field. The crowdsourcing tool is a web application that visualizes the test site on a 2D background map. The web application shows the pre-processed segment boundaries on the map and surveyors can annotate segments with the correct species and also add notes if needed. The crowdsourcing tool is presented in Fig. 4.
Figure 4: Screenshot of the crowdsourcing application used to collect ground-truth species data. In this example, the application shows a true orthophoto map, segment boundaries (red lines), the surveyor’s location (blue circle) based on GNSS positioning, and the information of the nearest tree together with possible actions for the surveyor. The true orthophoto map in the background was obtained from the City of Espoo (2024).
The crowdsourcing application enables surveyors to locate themselves on site with GNSS. To aid positioning when GNSS accuracy is low, a true orthophoto map from the City of Espoo (2024) was visualized as the background map. The 5-cm-resolution orthophoto map had been photographed in the summer of 2021. To help identifying the correct trees on site, the canopy height model of the test site was also added to the map as an additional map layer.
All segments were labeled by employees of the FGI. Most trees were labeled in the field. Some planted deciduous roadside trees were annotated using the open tree database of City of Espoo (2024), and some clear instances of spruces were labeled using the 5-cm-resolution true orthophoto due to the unique shape of the spruces.
Furthermore, the crowdsourcing tool was used as an on-site verification method for verifying species labels after the first labeling. For improved reliability of the dataset, we aimed to verify each labeled segment with another round of field visits. To help determine which segments needed more attention in the verification of the dataset, we computed for each segment the proportion of methods predicting an incorrect species and prepared an associated map. In the case of the more rare tree species, segments with multiple incorrect predictions were typically correctly labeled in the field. However, especially for the more common species, some segments with multiple incorrect predictions had been incorrectly labeled on the first round of field measurements, and they were subsequently corrected during the on-site verification.
@@ 2.6 Categories of tree segments
After labeling the segments and verifying the labels, the segments were categorized by manually inspecting profiles generated from the HeliALS segments. The applied categories were ‘Single tree’, ‘Large tree with undergrowth’, ‘Many trees of the same species’, ‘Many trees’, and ‘Tree section’. In the case of ‘Single trees’, there was only one tree visible in the profile image of the segment. A segment belongs to the ‘Large tree with undergrowth’ category if there was one taller tree and possibly some significantly smaller trees or bushes at most one third of the height of the larger tree in the segment. ‘Many trees of the same species’ category corresponds to segments which contained multiple trees, each having the same species. If there were multiple trees with different species in the same segment, the segment was categorized into the ‘Many trees’ category and the ground-truth species was recorded according to the species of the tallest tree within the segment. Segments with many trees where a single prominent tree species could not be identified were not included in the analysis. ‘Tree section’ corresponds to segments that contained only a part of a tree.
Figure 5 presents the profile category distribution by species for the test set. For each species class, the majority of the segments belong to the ‘Single tree’ category. However, it is worth mentioning that in the test site, alders typically grow in places with lots of other alders and, therefore, 30% of segments labeled as alder belong to the ‘Many trees of the same species’ category. Furthermore, 22% of maples and oaks belong to the ‘Tree section’ category because of over-segmentation of their wide canopies.
The segments were also categorized to different crown classes using alternative criteria based on tree heights and distances to neighboring trees as inspired by Wang et al. (2019a). Segments were considered ‘Isolated’ if within $8~{}\mathrm{m}$ radius, there were only segments with a height of at most half of the height of the ‘Isolated’ segment. Segments were considered ‘Dominant’ if the segment was not ‘Isolated’ and it was the tallest segment within $8~{}\mathrm{m}$ radius. If there was at least one segment at least $5~{}\mathrm{m}$ taller than a given segment within $6~{}\mathrm{m}$ radius, the segment was assigned to the ‘Smaller tree next to larger tree’ category. ‘Roadside’ segments were detected using the open tree database of the City of Espoo (2024). All other segments were assigned to the ‘Co-dominant’ category.
3 Algorithms for species classification
@@ 3.1 Overview of the benchmark contest
Participating partners were asked to submit the following: 1) a description of their applied methods, including relevant citations if previously published; 2) an Excel spreadsheet containing the tree segment number, classified species number, and test site code (A or B) for each segment (test site was divided into two areas); and 3) if modified segments (i.e., divided or merged) were used, a raster image of the final segment numbers, matching the original segment number raster in cell size and location.
@@ 3.2.1 Point-based (3D) deep learning methods
FGI-PointTransformer-DL-3D, HeliALS Removal of noise and ground points. Voxelization with a side length of 5 cm to obtain a more balanced point distribution. Point clouds of the three channels merged and each point assigned radiometric information from the two other channels based on nearest neighbor interpolation. (3+13)D point cloud with 8192 points per tree. Besides coordinates, attributes include intensities, amplitudes, reflectances, and echo deviations for the 3 scanners, and an echo return number. Coordinates normalized to unit sphere, other attributes normalized by scaling (except for the echo return number). Five randomly initialized Point Transformer models (Zhao et al., 2021) trained from scratch with 95-5 train-validation split. Both ordinary cross-entropy loss (FGI-PointTransformer-DL-3D) and weighted cross-entropy loss (FGI-PointTransformerWeighted-DL-3D) were tested. Augmentations applied at random (small random translation, scaling, and jittering of coordinates, and small random scaling and jittering of other attributes). Majority voting based on the five trained Point Transformer models.
@@
FGI-PointTransformer-VUX-DL-3D, HeliALS Same as FGI-PointTransformer-DL-3D but using radiometric and echo information only from the VUX scanner while ignoring radiometric information of the two other scanners. (3+6)D point cloud with 8192 points per tree. Besides point coordinates, attributes include the intensity, echo deviation, amplitude, reflectance, number of returns, and return number of the VUX scanner. Normalization as for FGI-PointTransformer-DL-3D. Same as FGI-PointTransformer-DL-3D. Both ordinary and weighted cross-entropy loss were tested. Same as FGI-PointTransformer-DL-3D.
@@
FGI-PointTransformer-VUXReflectance-DL-3D, HeliALS Same as FGI-PointTransformer-DL-3D but using only the reflectance from the VUX scanner and ignoring any other radiometric or echo information. (3+1)D point cloud with 8192 points per tree. Besides coordinates, attributes include the reflectance of the VUX scanner. Normalization as for FGI-PointTransformer-DL-3D. Same as FGI-PointTransformer-DL-3D. Both ordinary and weighted cross-entropy loss were tested. Same as FGI-PointTransformer-DL-3D.
@@
FGI-PointTransformer-GeometryOnly-DL-3D, HeliALS Same as FGI-PointTransformer-DL-3D but using only the geometry of the combined point cloud. 3D point cloud with 8192 points per tree. Coordinates normalized to unit sphere. Same as FGI-PointTransformer-DL-3D with ordinary cross-entropy loss. Same as FGI-PointTransformer-DL-3D.
@@
FGI-DGCNN-DL-3D, HeliALS Same as FGI-PointTransformer-DL-3D. Same as FGI-PointTransformer-DL-3D. Five randomly initialized dynamic graph convolutional neural net (DGCNN) models (Wang et al., 2019b) trained from scratch with 90-10 train-validation split and cross-entropy loss. Same augmentations as FGI-PointTransformer-DL-3D. Majority voting based on the five trained DGCNN models.
@@
FGI-PointNet-DL-3D, HeliALS Same as FGI-PointTransformer-DL-3D. Same as FGI-PointTransformer-DL-3D. Five randomly initialized PointNet models (Qi et al., 2017a) trained from scratch with 90-10 train-validation split and cross-entropy loss. Same augmentations as FGI-PointTransformer-DL-3D. Majority voting based on the five trained PointNet models.
@@
FGI-Point2Vec-DL-3D, HeliALS Same as FGI-PointTransformer-DL-3D, HeliALS. (3+13)D point cloud with 4096 points per tree. Besides coordinates, attributes include intensities, amplitudes, reflectances, and echo deviations for the 3 scanners, and an echo return number. Coordinates normalized to unit sphere, other attributes z-score normalized. Point2Vec model (Zeid et al., 2023) in fully-supervised classification setup. Trained from scratch using 80-20 train-validation split. Geometric augmentations applied at random (scaling, translation and rotation in $30^{\circ}$ degree steps around the z-axis). Weighted cross-entropy loss with label smoothing. Majority voting based on five inference runs with a single model.
@@
TUW-PointNet++-DL-3D, HeliALS Point clouds of the three channels combined by introducing three flag variables indicating the source channel of each point. (3+4)D point cloud with 8192 points per tree. Besides coordinates, attributes include the three flag variables (one of which is set to one), and the intensity value. Coordinates centered, but not scaled to preserve height. PointNet++ model (Qi et al., 2017b) with two PointNet layers and a final multilayer perceptron layer trained from scratch with 67-33 train-validation split. Geometric augmentations by generating 6 rotations around the z-axis for each tree. Weighted loss function to mitigate class imbalance. The model extends Ensemble PointNet++ (Winiwarter, 2024; Puliti et al., 2025) to multispectral data. Prediction of the trained model.
@@
FBK-PointNet++-DL-3D, HeliALS and Optech Titan Removal of noise and ground points. Point clouds of the three channels were merged and each point was assigned radiometric information from the three channels based on nearest neighbor interpolation. (3+3)D point cloud with 500 points per tree. Besides coordinates, attributes include the reflectance for each channel. Coordinates centered, but not scaled. Reflectances normalized. PointNet++ model (Qi et al., 2017b) trained from scratch with 80-20 train-validation split. Geometric and spectral augmentations for alder and oak to triple the number of examples. Geometric augmentations for rowan and linden to double the number of examples. Prediction of the trained model.
@@
FGI-PointTransformer-DL-3D, Optech Titan Removal of noise and ground points. Point clouds of the three channels merged and each point assigned intensity information from the two other channels based on nearest neighbors. (3+4)D point cloud with 2048 points per tree. Besides coordinates, attributes include intensities for the 3 scanners, and an echo return number. Normalization as for the HeliALS dataset. Same as FGI-PointTransformer-DL-3D for the HeliALS dataset using ordinary cross-entropy loss. Same as FGI-PointTransformer-DL-3D for the HeliALS dataset.
@@
In this work, our comparison includes the following 3D deep learning methods: FGI-PointTransformer-DL-3D and its variants, FGI-DGCNN-DL-3D, FGI-PointNet-DL-3D, FGI-Point2Vec-DL-3D, TUW-PointNet++-DL-3D, and FBK-PointNet++-DL-3D. The studied methods utilize various point-based deep learning models, including Point Transformer (Zhao et al., 2021) for FGI-PointTransformer-DL-3D, dynamic graph convolutional net (Wang et al., 2019b) for FGI-DGCNN-DL-3D, PointNet (Qi et al., 2017a) for FGI-PointNet-DL-3D, Point2Vec (Zeid et al., 2023) for FGI-Point2Vec-DL-3D, and PointNet++ (Qi et al., 2017b) for TUW-PointNet++-DL-3D and FBK-PointNet++-DL-3D. For most of the methods, nearest neighbor interpolation is used to represent the multispectral point cloud in a format, where each point contains the radiometric information from all of the three channels. As an exception, TUW-PointNet++-DL-3D does not perform nearest neighbor interpolation. Instead, each point is assigned a single intensity value, and three flag variables indicate the source channel of the intensity.
Importantly, there were differences in the used radiometric features and echo properties between the different models in the case of the HeliALS dataset that included the intensity, amplitude, reflectance, echo deviation and an echo return number for each point. FGI-PointTransformer-DL-3D, FGI-DGCNN-DL-3D, FGI-PointNet-DL-3D, and FGI-Point2Vec-DL-3D utilized all of the radiometric features and echo properties, implying that each 3D point was assigned 13 additional radiometric and echo features. TUW-PointNet++-DL-3D only utilized the intensities and FBK-PointNet++-DL-3D only considered the reflectances in the model input. For FGI-PointTransformer-DL-3D, we also studied different combinations of the included radiometric and echo features in order to evaluate their effect on the classification accuracy. In FGI-PointTransformer-VUX-DL-3D, we utilized the geometric information from the three scanners and radiometric and echo features only from the VUX scanner. As a further variant, FGI-PointTransformer-VUXReflectance-DL-3D only utilized the reflectance of the VUX scanner while disregarding all other radiometric and echo features. Finally, FGI-PointTransformer-GeometryOnly-DL-3D utilized no radiometric or echo information from any of the scanners, thus serving as a baseline when no radiometric information is available.
@@
Some of the studied methods attempt to mitigate the class imbalance of the training set. FGI-PointTransformerWeighted-DL-3D, FGI-Point2Vec-DL-3D and TUW-PointNet++-DL-3D utilize a weighted loss function to provide a higher weight to minority species, such as oak, alder and rowan. FBK-PointNet++-DL-3D instead augments the training dataset by generating geometric augmentations for rowan and linden, and geometric and spectral augmentations for alder and oak. On the other hand, TUW-PointNet++-DL-3D and the methods by FGI use augmentations to all segments during training without any specific focus on the minority species. The methods by FGI apply small random translation, scaling, and jittering of coordinates, as well as small random scaling and jittering of other attributes. TUW-PointNet++-DL-3D generates six rotational augmentations around the z-axis for each tree.

### v2

@@ 2.3 Pre-processing of laser scanning data
The data from each scanner were loaded to TerraScan (Terrasolid Ltd., Finland), and for each point the closest points within 20 cm were sought from the two other channels to form the reflectance triplets that were then stored in the RGB color fields in the LAZ file structure. Thus, points that do not have close-by counterparts on either one or two channels may exhibit only NA (coded as value ‘1’) on the respective RGB fields. The reflectance values in the LAZ were stored as decibel values and exported with RIEGL extrabytes (echo deviation, reflectance, and pulse width). The definition of the reflectance, amplitude and echo deviation attributes for the Riegl scanners are described in pfennigbauer2010improving. Eventually, the full point cloud dataset was split into 200 m square tiles and delivered for the benchmark in 1.2 LAZ format with ETRS-TM35FIN coordinates and ellipsoidal elevation values.
[…]
@@
Tree segmentation was carried out based on the Optech Titan data. The resulting segments were also applied for tree species classification with the HeliALS data to eliminate the variability of segment boundaries from the comparison. Given the fact that plenty of deciduous trees are growing in the study area, we used a varying window watershed-based method for the tree detection first depicted in (kaartinen2012international). First, we created a canopy height model (CHM) from the first returns of the vegetation points of Channels 1 and 2 with 0.5 m pixel size. The CHM was then smoothed by a Gaussian filter with a varying window size, followed by a local maximum filtering to identify the tops of trees. The window size varied according to the CHM height of the target as follows: 3 pixels for 0–7 m height, 5 pixels for 7–20 m height, 7 pixels for 20–30 m height, and 9 pixels for greater than 30 m height. Finally, the marker-controlled watershed algorithm was applied to delineate the crowns of trees using detected tree tops as seeds (kaartinen2012international). Output of the tree segmentation is a raster image with a unique label for each segment, i.e. the pixels with the same label form one segment. These images together with the original and normalized point clouds were delivered to the partners for species classification.
@@ 2.4 Benchmark dataset
Figure 3: (a) Example multispectral point clouds and photographs of the nine tree species found in the dataset. The red, green and blue color channels of the point clouds represent the laser pulse return intensities at the wavelengths $\lambda_{1}=\text{532 nm},\lambda_{2}=\text{905 nm}$ and $\lambda_{3}=\text{1550 nm}$, respectively. The height scale varies between the visualized point clouds. Photographs of the representative species instances were taken during the autumn of 2024. (b) Number of segments per species in the training and test sets. The dashed lines represent the training set (left y-axis) and the solid bars correspond to the test set (right y-axis). (c) Species-wise tree height distributions for the entire dataset.
The contestants of the benchmark competition received a training set containing segmented point clouds, the segment labels and the ground-truth species classes. The training set consisted of 1065 segments. The participating methods were benchmarked with a test set containing 5261 segments. The dataset was split into training and test sets using random sampling for a given number of segments per species. The full dataset was not available at the time of constructing the training set because the collection of the field reference data was still ongoing. Therefore, no detailed stratification approaches were used for splitting the dataset into training and test sets. Figure 3 visualizes the nine tree species included in the dataset, the number of segments per species in the training and test sets, and also the species-wise tree height distributions for the entire dataset. See also the distribution of mean intensity for the studied tree species at the three wavelengths of the HeliALS and Optech Titan systems in A. The distributions showcase the advantage of multispectral data in providing spectral information at multiple wavelengths.
@@ 2.5 Reference data collection
A browser-based crowdsourcing tool was developed to help collecting the ground-truth species data in the field. The crowdsourcing tool is a web application that visualizes the study area on a 2D background map as shown in Fig. 4. The web application shows the pre-processed segment boundaries on the map and surveyors can annotate segments with the correct species and also add notes if needed. The crowdsourcing application enables surveyors to locate themselves on the study area with GNSS. To aid positioning when GNSS accuracy is low, a true orthophoto map from the espoo_open_data was visualized as the background map. The 5-cm-resolution orthophoto map had been photographed in the summer of 2021. To help identifying the correct trees on the study area, the canopy height model of the study area was also added to the map as an additional map layer.
Figure 4: Screenshot of the crowdsourcing application used to collect ground-truth species data. In this example, the application shows a true orthophoto map, segment boundaries (red lines), the surveyor’s location (blue circle) based on GNSS positioning, and the information of the nearest segment together with possible actions for the surveyor. The true orthophoto map in the background was obtained from the espoo_open_data.
Reference data was collected in the field using the crowdsourcing application by the employees of the FGI in 2023 and 2024. The process was guided by several key principles. First, participants were introduced to the data collection tools before starting the field work to establish a baseline for consistency and reduce initial misclassification errors. Second, we made improvements to the user interface of the crowdsourcing application during the field campaign based on our learnings in the field to further ease the collection of the ground-truth data. Third, we implemented an expert verification process to ensure data quality and reduce inherent subjectivity: The initial species classifications made by the crowd were subsequently revisited on the field and verified by experts using the crowdsourcing tool. To help determine which segments needed more attention in the verification phase, we computed for each segment the proportion of classification methods that predicted an incorrect species and prepared an associated map. In the case of minority species, segments with multiple incorrect predictions were typically correctly labeled in the field. For the more common species, some segments with multiple incorrect predictions had been incorrectly labeled on the first round of field measurements, and they were subsequently corrected during the on-site verification. This two-step approach—using the crowd for broad coverage and experts for final validation—helped to reduce subjectivity.
In addition to field data collection using the crowdsourcing application, some planted deciduous roadside trees were annotated using the open tree database of espoo_open_data, and some clear instances of spruces were labeled using the 5-cm-resolution true orthophoto due to the unique shape of the spruces.
@@ 2.6 Categories of tree segments
After labeling the segments and verifying the labels, the segments were categorized by manually inspecting profiles generated from the HeliALS segments. The applied categories were ‘Single tree’, ‘Large tree with undergrowth’, ‘Many trees of the same species’, ‘Many trees of many species’, and ‘Tree section’. In the case of ‘Single trees’, there was only one tree visible in the profile image of the segment. A segment belongs to the ‘Large tree with undergrowth’ category if there was one taller tree and possibly some significantly smaller trees or bushes at most one third of the height of the larger tree in the segment. ‘Many trees of the same species’ category corresponds to segments which contained multiple trees, each having the same species. If there were multiple trees with different species in the same segment, the segment was categorized into the ‘Many trees of many species’ category and the ground-truth species was recorded according to the species of the tallest tree within the segment. Segments with many trees where a single prominent tree species could not be identified were not included in the analysis. ‘Tree section’ corresponds to segments that contained only a part of a tree.
Figure 5 presents the profile category distribution by species for the test set. For each species class, the majority of the segments belong to the ‘Single tree’ category. However, it is worth mentioning that in the study area, alders typically grow in places with lots of other alders and, therefore, 30% of segments labeled as alder belong to the ‘Many trees of the same species’ category. Furthermore, 22% of maples and oaks belong to the ‘Tree section’ category because of over-segmentation of their wide canopies.
The segments were also categorized to different crown classes using alternative criteria based on tree heights and distances to neighboring trees as inspired by wang2019field. Segments were considered ‘Isolated’ if within $8~\mathrm{m}$ radius, there were only segments with a height of at most half of the height of the ‘Isolated’ segment. Segments were considered ‘Dominant’ if the segment was not ‘Isolated’ and it was the tallest segment within $8~\mathrm{m}$ radius. If there was at least one segment at least $5~\mathrm{m}$ taller than a given segment within $6~\mathrm{m}$ radius, the segment was assigned to the ‘Smaller tree next to larger tree’ category. ‘Roadside’ segments were detected using the open tree database of the espoo_open_data. All other segments were assigned to the ‘Co-dominant’ category.
3 Benchmark contest and species classification algorithms
@@ 3.1 Overview of the benchmark contest
Participating partners were asked to submit the following: 1) a description of their applied methods, including relevant citations if previously published; 2) an Excel spreadsheet containing the tree segment number, classified species number, and sub-region code (A or B) for each segment (study area was divided into two sub-regions); and 3) if modified segments (i.e., divided or merged) were used, a raster image of the final segment numbers, matching the original segment number raster in cell size and location.
@@ 3.2.1 Point-based (3D) deep learning methods
FGI-PointTransformer-DL-3D, HeliALS Removal of noise and ground points. Voxelization with a side length of 5 cm to obtain a more balanced point distribution. Point clouds of the three channels merged and each point assigned radiometric information from the two other channels based on nearest neighbor interpolation. Imputation strategies used if point attributes are missing for some of the channels. (3+13)D point cloud with 8192 points per segment. Besides coordinates, attributes include intensities, amplitudes, reflectances, and echo deviations for the 3 scanners, and an echo return number. Coordinates normalized to unit sphere, other attributes normalized by scaling (except for the echo return number). Five randomly initialized Point Transformer models (zhao2021point) trained from scratch with 95-5 train-validation split. Both ordinary cross-entropy loss (FGI-PointTransformer-DL-3D) and weighted cross-entropy loss (FGI-PointTransformerWeighted-DL-3D) were tested. Augmentations applied at random (small random translation, scaling, and jittering of coordinates, and small random scaling and jittering of other attributes). Majority voting based on the five trained Point Transformer models.
@@
FGI-PointTransformer-VUX-DL-3D, HeliALS Same as FGI-PointTransformer-DL-3D but using radiometric and echo information only from the VUX scanner while ignoring radiometric information of the two other scanners. (3+6)D point cloud with 8192 points per segment. Besides point coordinates, attributes include the intensity, echo deviation, amplitude, reflectance, number of returns, and return number of the VUX scanner. Normalization as for FGI-PointTransformer-DL-3D. Same as FGI-PointTransformer-DL-3D. Both ordinary and weighted cross-entropy loss were tested. Same as FGI-PointTransformer-DL-3D.
@@
FGI-PointTransformer-VUXReflectance-DL-3D, HeliALS Same as FGI-PointTransformer-DL-3D but using only the reflectance from the VUX scanner and ignoring any other radiometric or echo information. (3+1)D point cloud with 8192 points per segment. Besides coordinates, attributes include the reflectance of the VUX scanner. Normalization as for FGI-PointTransformer-DL-3D. Same as FGI-PointTransformer-DL-3D. Both ordinary and weighted cross-entropy loss were tested. Same as FGI-PointTransformer-DL-3D.
@@
FGI-PointTransformer-GeometryOnly-DL-3D, HeliALS Same as FGI-PointTransformer-DL-3D but using only the geometry of the combined point cloud. 3D point cloud with 8192 points per segment. Coordinates normalized to unit sphere. Same as FGI-PointTransformer-DL-3D with ordinary cross-entropy loss. Same as FGI-PointTransformer-DL-3D.
@@
FGI-DGCNN-DL-3D, HeliALS Same as FGI-PointTransformer-DL-3D. Same as FGI-PointTransformer-DL-3D. Five randomly initialized dynamic graph convolutional neural net (DGCNN) models (wang2019dynamic) trained from scratch with 90-10 train-validation split and cross-entropy loss. Same augmentations as FGI-PointTransformer-DL-3D. Majority voting based on the five trained DGCNN models.
@@
FGI-PointNet-DL-3D, HeliALS Same as FGI-PointTransformer-DL-3D. Same as FGI-PointTransformer-DL-3D. Five randomly initialized PointNet models (qi2017pointnet) trained from scratch with 90-10 train-validation split and cross-entropy loss. Same augmentations as FGI-PointTransformer-DL-3D. Majority voting based on the five trained PointNet models.
@@
FGI-Point2Vec-DL-3D, HeliALS Same as FGI-PointTransformer-DL-3D, HeliALS. (3+13)D point cloud with 4096 points per segment. Besides coordinates, attributes include intensities, amplitudes, reflectances, and echo deviations for the 3 scanners, and an echo return number. Coordinates normalized to unit sphere, other attributes z-score normalized. Point2Vec model (zeid2023point2vec) in fully-supervised classification setup. Trained from scratch using 80-20 train-validation split. Geometric augmentations applied at random (scaling, translation and rotation in $30^{\circ}$ degree steps around the z-axis). Weighted cross-entropy loss with label smoothing. Majority voting based on five inference runs with a single model.
@@
TUW-PointNet++-DL-3D, HeliALS Point clouds of the three channels combined by introducing three flag variables indicating the source channel of each point. (3+4)D point cloud with 8192 points per segment. Besides coordinates, attributes include the three flag variables (one of which is set to one), and the intensity value. Coordinates centered, but not scaled to preserve height. PointNet++ model (qi2017pointnet++) with two PointNet layers and a final multilayer perceptron layer trained from scratch with 67-33 train-validation split. Geometric augmentations by generating 6 rotations around the z-axis for each segment. Weighted loss function to mitigate class imbalance. The model extends Ensemble PointNet++ (TUWien_github2024; puliti2025benchmarking) to multispectral data. Prediction of the trained model.
@@
FBK-PointNet++-DL-3D, HeliALS and Optech Titan Removal of noise and ground points. Point clouds of the three channels were merged and each point was assigned radiometric information from the three channels based on nearest neighbor interpolation. (3+3)D point cloud with 500 points per segment. Besides coordinates, attributes include the reflectance for each channel. Coordinates centered, but not scaled. Reflectances normalized. PointNet++ model (qi2017pointnet++) trained from scratch with 80-20 train-validation split. Geometric and spectral augmentations for alder and oak to triple the number of examples. Geometric augmentations for rowan and linden to double the number of examples. Prediction of the trained model.
@@
FGI-PointTransformer-DL-3D, Optech Titan Removal of noise and ground points. Point clouds of the three channels merged and each point assigned intensity information from the two other channels based on nearest neighbors. (3+4)D point cloud with 2048 points per segment. Besides coordinates, attributes include intensities for the 3 scanners, and an echo return number. Normalization as for the HeliALS dataset. Same as FGI-PointTransformer-DL-3D for the HeliALS dataset using ordinary cross-entropy loss. Same as FGI-PointTransformer-DL-3D for the HeliALS dataset.
@@
In this work, our comparison includes the following 3D deep learning methods: FGI-PointTransformer-DL-3D and its variants, FGI-DGCNN-DL-3D, FGI-PointNet-DL-3D, FGI-Point2Vec-DL-3D, TUW-PointNet++-DL-3D, and FBK-PointNet++-DL-3D. The studied methods utilize various point-based deep learning models, including Point Transformer (zhao2021point) for FGI-PointTransformer-DL-3D, dynamic graph convolutional net (wang2019dynamic) for FGI-DGCNN-DL-3D, PointNet (qi2017pointnet) for FGI-PointNet-DL-3D, Point2Vec (zeid2023point2vec) for FGI-Point2Vec-DL-3D, and PointNet++ (qi2017pointnet++) for TUW-PointNet++-DL-3D and FBK-PointNet++-DL-3D. For most of the methods, nearest neighbor interpolation is used to represent the multispectral point cloud in a format, where each point contains the radiometric information from all of the three channels. As an exception, TUW-PointNet++-DL-3D does not perform nearest neighbor interpolation. Instead, each point is assigned a single intensity value, and three flag variables indicate the source channel of the intensity.
There were differences in the used radiometric features and echo properties between the different models in the case of the HeliALS dataset that included the intensity, amplitude, reflectance, echo deviation and an echo return number for each point. FGI-PointTransformer-DL-3D, FGI-DGCNN-DL-3D, FGI-PointNet-DL-3D, and FGI-Point2Vec-DL-3D utilized all of the radiometric features and echo properties, implying that each 3D point was assigned 13 additional radiometric and echo features. TUW-PointNet++-DL-3D only utilized the intensities and FBK-PointNet++-DL-3D only considered the reflectances in the model input. For FGI-PointTransformer-DL-3D, we also studied different combinations of the included radiometric and echo features in order to evaluate their effect on the classification accuracy. In FGI-PointTransformer-VUX-DL-3D, we utilized the geometric information from the three scanners and radiometric and echo features only from the VUX scanner. As a further variant, FGI-PointTransformer-VUXReflectance-DL-3D only utilized the reflectance of the VUX scanner while disregarding all other radiometric and echo features. Finally, FGI-PointTransformer-GeometryOnly-DL-3D utilized no radiometric or echo information from any of the scanners, thus serving as a baseline when no radiometric information is available.
@@
Some of the studied methods attempt to mitigate the class imbalance of the training set. FGI-PointTransformerWeighted-DL-3D, FGI-Point2Vec-DL-3D and TUW-PointNet++-DL-3D utilize a weighted loss function to provide a higher weight to minority species, such as oak, alder and rowan. FBK-PointNet++-DL-3D instead augments the training dataset by generating geometric augmentations for rowan and linden, and geometric and spectral augmentations for alder and oak. On the other hand, TUW-PointNet++-DL-3D and the methods by FGI use augmentations to all segments during training without any specific focus on the minority species. The methods by FGI apply small random translation, scaling, and jittering of coordinates, as well as small random scaling and jittering of other attributes. TUW-PointNet++-DL-3D generates six rotational augmentations around the z-axis for each segment.